OpenStudy (anonymous):

Currently, Sam's coach is twice as old as Sam. When Sam started soccer 10 years ago, his coach was four times his age. How old is Sam now?

OpenStudy (anonymous):

2s = c 4(s-10) = c-10 //set up equations 4s - 40 = c -10 //simplify equations 4s - 40 + 10 = c //get c by itself 4s - 30 = c //substitute 2s = 4s - 30 //simplify 30 = 4s - 2s 30 = 2s s = 15 Same is 15 //check sam is 15 coach is 30 10 years ago sam is 5 coach is 20